find the slope of 3x+y=4, finding the slope of things like that is really easy.... try to always change things into y=mx+b. y is y and the number before x is your slope and b is the whole number with no x or y....
Answer by smallfry(7)

finding the slope of things like that is really easy.... try to always change things into y=mx+b. y is y and the number before x is your slope and b is the whole number with no x or y....
so....
3x+y=4
substract 3x to get y alone..
3x=y=4
-3x -3x ( do it to both sides)
y=4-3x...
since -3 is before x.. thats your slope to graph that it would be 3/1.
** make sure y is always only one y...